### v1

# Mining Legal Arguments to Study Judicial Formalism

Abstract
Courts must justify their decisions, but systematically analyzing judicial reasoning at scale remains difficult. This study refutes claims about formalistic judging in Central and Eastern Europe (CEE) by developing automated methods to detect and classify judicial reasoning in Czech Supreme Courts’ decisions using state-of-the-art natural language processing methods. We create the MADON dataset of 272 decisions from two Czech Supreme Courts with expert annotations of 9,183 paragraphs with eight argument types and holistic formalism labels for supervised training and evaluation. Using a corpus of 300k Czech court decisions, we adapt transformer LLMs for Czech legal domain by continued pretraining and experiment with methods to address dataset imbalance including asymmetric loss and class weighting. The best models successfully detect argumentative paragraphs (82.6% macro-F1), classify traditional types of legal argument (77.5% macro-F1), and classify decisions as formalistic/non-formalistic (83.2% macro-F1). Our three-stage pipeline combining ModernBERT, Llama 3.1, and traditional feature-based machine learning achieves promising results for decision classification while reducing computational costs and increasing explainability. Empirically, we challenge prevailing narratives about CEE formalism. This work shows that legal argument mining enables reliable judicial philosophy classification and shows the potential of legal argument mining for other important tasks in computational legal studies. Our methodology is easily replicable across jurisdictions, and our entire pipeline, datasets, guidelines, models, and source codes are available at https://github.com/trusthlt/madon/.
keywords:
Legal argument mining, judicial formalism, large language models, computational legal studies, Central and Eastern Europe
1 Introduction
Courts are expected to justify their decisions. How courts reason matters to scholars, practitioners, and citizens alike [Walton_Macagno_Sartor_2021]. Yet systematically analyzing judicial reasoning at scale remains difficult. It is a challenging interdisciplinary task that inherits the key problems from legal NLP, legal theory, and empirical legal studies. Empirical work often depends on manual annotations, which are time-consuming and costly [hall2008, gilardi_chatgpt_2023, braun_i_2023-1]. Legal concepts like “formalism” or traditional argument types remain disputed or abstract, hard to measure across large numbers of cases and foreign to NLP community [luders_classifying_2024-1, habernal_mining_2024-1, schauer_1988]. While recent advances in NLP and legal argument mining promise to support quantitative computational analysis of legal reasoning, key challenges persist: clarifying complex legal concepts [habernal_mining_2024-1, luders_classifying_2024-1], addressing scarce, imbalanced annotations with different quality [braun_i_2023-1], handling non-English legal texts, and ensuring explainable, reliable pipelines [Chau_Livermore_2024]. In short, judicial reasoning matters, but studying it at scale remains hard.
We turn to Central and Eastern Europe (CEE), a region rarely examined empirically, yet widely alleged to be formalistic. Scholarship often portrays CEE judiciaries as rooted in late-communist legal culture and narrowly focused on text and procedure [cserne_discourses_2020-1, kuhn_judiciary_2011-2, matczak_eu_2015, matczak_constitutions_2010-2, manko_weeds_2013, bystranowski_formalist_2022-1, Bencze_2021], but much of this is anecdotal. We still know little about how CEE courts in fact justify their decisions [cserne_discourses_2020-1, cserne_judicial_2024, komarek_2015], and, to our knowledge, no study has used LLMs to test whether formalism indeed prevails in the region. Czechia offers an ideal case study: the Supreme Court (SC), largely inherited from the communist era, and the Supreme Administrative Court (SAC), founded in 2003 and seen as more reformist, are routinely described as opposites—rigidly formalistic SC versus more discursive SAC [matczak_eu_2015, kosar_domestic_2020-1, sipulova_kosar_purging_2024-1, kuhn_worlds_2004, kuhn_application_2005, kuhn_judiciary_2011-2, kuhn_quality_2018, matczak_constitutions_2010-2, Kadlec2016, stehlik_unijni_2014]. We use this setting to test the broader claims about allegedly omnipresent formalism in CEE as well as to evaluate the potential of legal NLP on non-English judgments.
We ask three research questions, namely 1) what are the capabilities of state-of-the-art NLP methods to detect and classify traditional types of legal arguments in non-english judicial decisions, 2) can they distinguish between formalistic and non-formalistic reasoning styles at the decision level, and 3) does empirical analysis support or refute prevailing claims about CEE courts?
By addressing these research questions, we make five substantial contributions to the field. First, we create the MADON dataset: a collection of 272 Czech Supreme Court and Supreme Administrative Court decisions with ca 9,183 paragraphs anno- tated by legal experts with eight argument types and one holistic formalism label. Second, we publish detailed annotation guidelines grounded in established theories of legal argumentation and refined through over 200 pilot cases. Third, we compile and preprocess a corpus of 300,000 anonymized Czech court decisions for continued pretraining of LLMs. Fourth, we experiment with pretraining, fine-tuning and testing different configurations of state-of-the-art open models (Llama 3.1, ModernBERT) across three tasks: detecting arguments, classifying argument types, and predicting formalism. From these results, we propose a unified pipeline to classify decisions as non/formalistic that integrates all tasks and model families, improving explainability, reducing computational costs, and boosting performance. Finally, we use the MADON dataset for first empirical analysis of the two Czech apex courts. In sum, we show how contemporary NLP research helps to investigate judicial philosophies and traditional jurisprudential questions; we provide novel data, pipelines, datasets and models for such investigations.
2 Related work
2.1 Jurisprudential debate on formalism
A simple example illustrates formalistic judging. Suppose a statute fixes a 15-day deadline to file an appeal. If a party files on day 16 because the counsel was hospitalized, a formalistic judge would dismiss the appeal as inadmissible, since the statute contains no exception. A non-formalistic judge would accept the appeal, invoking access-to-justice principles, the purpose of the rule, and the unfairness of harsh consequences under the individual circumstances.
Formalism concerns how lawyers, typically judges, approach legal questions. It is an approach to law and judicial adjudication that prioritizes the binding authority of legal norms and their wording. A formalist sees law as distinct from politics and limits the influence of non-legal norms and extra-legal considerations in answering legal questions, such as social, economical, political, or moral factors [weber_1922, schauer_1988, Weinrib1988].
The jurisprudential debate on legal formalism is longstanding and widespread. Some authors argue that the term has lost all meaning, while others have sought to rediscover and clarify it [schauer_1988, Nachbar2020]. The debate spans sociology of law [weber_1922], jurisprudence and legal argumentation [schauer_1988], Law and Economics [Djankov_2002] and critical legal studies [Kennedy2004Disenchantment]. Despite the variety of perspectives on formalism, prominent scholars from these different traditions usually discuss the question already noted at the outset: to what extent are the judges bound by existing legal rules and their wording?
Formalism can be approached analytically, descriptively, or normatively. The analytical approach asks, e.g., what formalism is and whether it is possible (e.g. can a judge simply apply a rule given the very same legal system already contains open-textured principles). The descriptive approach asks whether judges in fact decide cases in a formalistic manner (e.g. would an average judge indeed reject the appeal on strictly procedural grounds?). The normative approach asks whether formalism is desirable and what judges should do when deciding cases - is it right that judges stick to statutory wording rather than purpose?
2.2 Legal formalism in Central and Eastern Europe (CEE)
Both descriptive and normative approaches toward formalism are especially prominent in the CEE.
Descriptively, formalism is widely portrayed as the dominant judicial style in the CEE region [bystranowski_formalist_2022-1, dixon_2023, Bencze_2021]. Existing scholarship traces the formalism to socialist-era legal culture and unreformed post-communist judiciary [cserne_discourses_2020-1, kosar_comparative_2023, manko_weeds_2013, Suteu_2023]. Labels include “hyperpositivism” [manko_weeds_2013], “mechanical jurisprudence” [kosar_comparative_2023], and “vulgar formalism” [Suteu_2023]. This portrayal appears across the region, in Poland [Kustra-Rogatka_2023, manko_weeds_2013], Czechia [kuhn_judiciary_2011-2, kosar_domestic_2020-1, Jakab_Dyevre_Itzcovich_2017], Romania [Suteu_2023], Serbia [Besirevic_2014], and parts of the post-Yugoslav space [uzelac_survival_2010]. It leads some theorists to endorse very strong claims suggesting that CEE legal systems might collectively form a distinct ”third legal family” alongside common and civil law traditions [manko_2016]. All this because the CEE judiciary is allegedly formalistic.
From the normative perspective, formalism is heavily criticized in the CEE region. Scholars argue that it impedes rule of law [matczak_2018] and the enforcement of EU law [matczak_constitutions_2010-2, kuhn_worlds_2004]. According to kuhn_judiciary_2011-2, the formalism shows that the transition to liberal democracy was weak or incomplete, given the alleged communist-era roots of formalism. These criticisms are echoed in practice. The Czech Constitutional Court has repeatedly criticized the rest of the judiciary for excessive formalism and struck excessive formalism down as a violation of fundamental rights [constitutionalcourt2012]. Alleged formalism has also been raised in judicial appointments to constitutional courts as a reason to nominate more liberal judges [CzechSenate2002]. Beyond the scholarship and courts, politicians have critiqued formalism too: the Slovak President recently blamed judicial formalism for public frustration in post-communist Slovakia and Czechia [CT24_2024_Caputova]. Formalism has become central to debates about the region’s communist past and democratic future. Key actors view formalism as a toxic post-communist legacy undermining judicial credibility and the rule of law in the new democratic state.
The formalism debate resonates particularly strongly in Czechia. There, it often concerns the country’s two supreme courts: the Supreme Court (SC), which handles civil and criminal matters, and the Supreme Administrative Court (SAC), deciding administrative law matters. The SC is characterized as an unreformed formalistic institution tied to the communist past, similar to most other courts in the judicial system [kuhn_worlds_2004, kuhn_application_2005, kuhn_judiciary_2011-2, kuhn_quality_2018, matczak_constitutions_2010-2, matczak_eu_2015, kosar_domestic_2020-1, sipulova_kosar_purging_2024-1, Kadlec2016, stehlik_unijni_2014]. Indeed, the SC was “inherited” from the communist era and staffed with many older judges who were formerly members of the communist party [sipulova_kosar_purging_2024-1]. By contrast, the newly established Supreme Administrative Court (SAC), created in 2003 to adjudicate administrative matters and reportedly staffed by new generation of jurists, is described as “discursive” and more open in its reasoning [matczak_eu_2015]. For over two decades, scholars have raised the distinct genealogies of these two courts as evidence of their divergent decision-making styles [kuhn_worlds_2004, kuhn_application_2005, kuhn_judiciary_2011-2, kuhn_quality_2018, matczak_constitutions_2010-2, matczak_eu_2015, kosar_domestic_2020-1, sipulova_kosar_purging_2024-1, Kadlec2016, stehlik_unijni_2014]. We call this the Tale of Two Courts.
Despite the omnipresent debates about formalism, however, little is known about how CEE judiciaries actually decide cases. One trend in the CEE literature [komarek_2015, cserne_judicial_2024, manko_weeds_2013] rightly emphasizes that claims of pervasive formalism are mostly anecdotal and based on the views of a few elite insiders, rather than systematic evidence [cserne_discourses_2020-1]. Without reliable data about judicial reasoning in the region, prevailing claims about CEE courts remain unsubstantiated. This applies equally to the Tale of Two Courts itself as there is no empirical study on the adjudication of SC and its reasoning. As [cserne_judicial_2024] concludes, empirical analysis of formalism in CEE requires clear definitions of formalism and remains in its early stages. Simply put, the evidence is missing.
This paper aims to close the existing research gap by a pioneering empirical research in this direction.
2.3 Empirical analysis of formalism
We now discuss existing studies that investigate formalism empirically—both within the CEE region and beyond.
Three main approaches can be distinguished. The first approach studies the types of arguments courts employ. It quantifies the frequency of various argument types (e.g., linguistic, systematic, teleological) that are then classified as either formalistic (e.g., linguistic arguments) or non-formalistic (e.g., teleological arguments). This approach is not new; legal theory has investigated argument types for centuries, and typically US scholars like choi_2020 or Krishnakumar2020, krishnakumar_textualism_2024 have been empirically investigating argument types to assess prevailing interpretative theories. Fewer studies focus on argument types to measure formalism specifically. Most of these concern CEE judiciaries [matczak_constitutions_2010-2, matczak_eu_2015, bystranowski_formalist_2022-1]. The underlying assumption of this approach is that the more formalistic arguments a court uses, the more formalistic the court is.
The second approach [stiglitz_h_historical_2024, thalken_modeling_2023] directly labels paragraphs as formalistic or non-formalistic without analyzing specific argument types. This simplification enabled Stiglitz and Thalken not only to manually code decisions but also to provide large-scale analysis using NLP. This approach assumes that the higher the number of formalistic paragraphs appearing in courts’ decisions, the more formalistic the court is.
The third approach captures formalism’s different dimensions by examining various parameters simultaneously, including the court’s rhetorical style and procedural choices alongside argumentation patterns. This approach offers a holistic perspective on how formalism manifests across different dimensions of judicial decision-making [alberstein_sarat_measuring_2012, alberstein_whats_2020].
Despite valuable contributions, current approaches to measuring legal formalism face several problems. Most rely on small, manually coded samples without NLP, and none use argument mining to measure formalism directly. They typically cover only narrow slices of case law—e.g., a handful of criminal cases [bystranowski_formalist_2022-1] or selected administrative cases [matczak_constitutions_2010-2, matczak_eu_2015]. More importantly, many suffer from methodological shortcomings: missing or insufficient coding manuals and a lack of reported inter-annotator agreement, which undermines reproducibility [matczak_eu_2015, bystranowski_formalist_2022-1].
Each of the three approaches has additional limitations. First, the argument type approach can be insufficient: simple counts often miss crucial nuances. While this method yields useful quantitative data about which argument types courts use and how often, it can mislead as a measure of formalism when, for example, non-formalistic arguments are typically decisive but scarce while formalistic ones appear very often but are merely supplementary or peripheral. Simply put, a single well-put policy argument may outweigh multiple brief references to the wording, when assessing formalistic nature of a decision.
Second, the paragraph labeling approach likewise downplays the importance of particular passages. It also downplays argument types, the standard lens in legal theory and legal argumentation [Walton_Macagno_Sartor_2021, alexy_argument_2010] and provides less informative data on reasoning practices.
Finally, the multiparameter approach, though comprehensive, brings significant complexity to annotation.
We differentiate substantially from previous works and combine the best of all worlds. First, we analyze argument types and go further by publishing detailed annotation guidelines, reporting inter-annotator agreement, and employing state of the art LLMs. Second, we combine fine-grained argument mining classifying paragraphs into 8 argument types with a binary decision level label formalistic/non-formalistic; to this end, we design unique pipeline that reduces computational requirements, increases explainability and boosts performance. Third, we consider multiple aspects of formalism as the last approach, but condense them into a single binary holistic label, formalistic or non formalistic, and also employ NLP to predict this variable. As such, our approach builds on existing methods, but enables more comprehensive measurement of judicial reasoning and formalism.
The discussed shortcomings of existing approaches show the need for methods that remain true to legal theory but can work on a larger scale. Legal argument mining offers such a path.
2.4 Legal argument mining
Argument mining aims to automatically identify and extract arguments [lawrence_argument_2019]. Legal argument mining (LAM) focuses on the legal domain and legal arguments [habernal_mining_2024-1, zhang2022, MochalesMoens2011]. Recent research has applied legal argument mining to classify the traditional types of legal arguments [habernal_mining_2024-1, luders_classifying_2024-1]. This emerging field has been growing in recent years and increasingly intersects with empirical legal research. For recent advances, we refer to recent LAM papers [habernal_mining_2024-1, luders_classifying_2024-1] as well as a systematic literature review of the field [zhang2022].
Several related studies, while not framed as LAM, also classify legal arguments or interpretive theories in judgments using machine learning and NLP. Examples include Peters2024 on US state courts, DugacAltwicker2025 on the ECtHR, stiglitz_h_historical_2024 on the entire U.S. Supreme Court corpus (1870–2023), MunozSoro_etal_2024 on Spanish custody judgments or work on U.S. tax statutory interpretation using traditional ML techniques [choi_2020].
To our knowledge, however, no study has used LLMs or other ML methods to examine interpretive theories of CEE courts.
In sum, despite a century of debate about formalism and decades of strong claims about CEE judicial formalism, the field lacks both clear definitions and systematic evidence. Previous empirical studies remain small-scale and unreproducible, while legal argument mining has yet to be applied to CEE courts. We fill this gap by combining prominent works in legal theory with state-of-the-art NLP methods to test these longstanding claims at scale, as the next section explains.
3 Measuring formalism in judicial decisions
3.1 Five core tenets of CEE formalism
Current legal scholarship on formalism in CEE converges on five distinctive criteria of formalistic judicial decision-making. Formalistic courts in the CEE would
  1. 1.
    rely heavily on a limited set of arguments derived from statutory text [matczak_constitutions_2010-2, matczak_eu_2015, bystranowski_formalist_2022-1, Jakab_Dyevre_Itzcovich_2017, cserne_discourses_2020-1, kuhn_judiciary_2011-2],
  2. 2.
    focus on the most locally applicable rule, not broader principles [matczak_constitutions_2010-2, matczak_eu_2015, bystranowski_formalist_2022-1, manko_weeds_2013, cserne_discourses_2020-1, Bencze_2021, kuhn_judiciary_2011-2],
  3. 3.
    usually exclude “external standards” in reasoning, like efficiency, justice, moral and political reasoning [kuhn_worlds_2004, matczak_constitutions_2010-2, bystranowski_formalist_2022-1, MalolepszyGluchowski2021], teleological interpretation [kuhn_worlds_2004, kuhn_judiciary_2011-2, Jakab_Dyevre_Itzcovich_2017, MalolepszyGluchowski2021] or even historical interpretation [matczak_constitutions_2010-2, matczak_eu_2015],
  4. 4.
    dismiss cases on formal grounds to avoid analyzing them on the merits [manko_weeds_2013, bystranowski_formalist_2022-1, uzelac_survival_2010], and
  5. 5.
    provide scarce reasoning for their decisions [kuhn_worlds_2004, Suteu_2023, kuhn_judiciary_2011-2].
We call these five criteria the core tenets of CEE formalism.
3.2 Measuring formalism through argument types
The key literature on formalism discusses legal argumentation; the three core tenets of CEE formalism concern arguments. We therefore begun with arguments.
On their own, the tenets are too abstract for empirical coding of concrete cases. For example, the first core tenet of CEE formalism (heavy reliance on text based arguments) requires a precise specification: what counts as a text-based argument, and how can we identify the linguistic argument is really present in an actual decision? Without such specification, empirically analyzing real world case law is challenging, if not impossible.
To address this challenge, we developed a comprehensive annotation guidelines including an annotation scheme and instructions specifying what each type means and when to annotate it. The guidelines emerged iteratively: we began with leading argumentation taxonomies [alexy_argument_2010, maccormick_interpreting_2016, Walton_Macagno_Sartor_2021], incorporated German and Czech legal argumentation scholarship to represent literature from the Cetral European region [Mollers2020LegalMethods, Wintr2019Metody, LarenzCanaris_Methodenlehre_1995], applied and refined them on more than 200 Czech Supreme Court pilot decisions (SC and SAC), and then re-checked the adjustments against the same literature. We repeated these steps iteratively over several months before annotating the MADON dataset.
Our guidelines identified eight argument types, which we classify as formalistic or non-formalistic based on their relationship to the core tenets of CEE formalism:
Formalistic argument types
LIN
– Linguistic Interpretation
SI
– Systemic Interpretation (including Constitutional‑conforming and EU‑conforming interpretation)
CL
– Case Law
D
– Doctrine
Non-formalistic argument types
HI
– Historical Interpretation
PL
– Principles of Law and Values (including Constitutional Values, Rights, Principles and EU Values and Principles)
TI
– Teleological Interpretation
PC
– Practical Consequences
For each argument type, our guideliness included a structured description that connects the theoretical concept to real‑world decisions. For most types, we (1) give a brief description, (2) list its subtypes, (3) note “trigger” phrases that usually signal its presence in a decision, (4) provide typical and borderline examples, (5) add comments and annotator tips and (6) include annotation chart consistent with [thalken_modeling_2023] to guide the annotation process. The creation of these guidelines itself represents an important contribution to the field. [^1] Appendix A contains the summary of the guideliness. Full version of the guideliness is accesible at https://github.com/trusthlt/madon/.
We then applied the guideliness to annotate the below described MADON dataset of 272 decisions from Supreme Court and Supreme Administrative Court of Czechia. We used set of cases distinct from the pilot decisions to support reliability of the annotation.
The court decisions are structured into paragraphs. We reflected this structure and chose paragraph-level annotation over sentence-level coding for annotating arguments. Sentence-level coding produced very low Krippendorff unitized Alpha scores—annotators disagreed where exactly arguments start and end. From the perspective of our research question, we were not so much interested in whether a particular sentence belonged to an argument, but more about whether a particular argument is present in a certain paragraph. Thus, this decision should not negatively influence the results.
Each argument type would be counted only once per paragraph, even if multiple instances within one paragraph appeared. However, one paragraph could include two different argument types. For example, if a paragraph cited five previous judgments and referenced two explanatory notes, we would code it once for case law (CL – Case Law) and once for historical interpretation (HI – Historical Interpretation). We hypothesize that this approach reflects how legal theorists typically analyze argumentation: focusing on argument types rather than counting individual citations within one single sentence or paragraph.
We excluded rejected arguments from annotation. When courts discussed competing interpretations, we coded only the reasoning ultimately adopted. For example, if a court noted that historical interpretation supported outcome X while linguistic interpretation pointed to non-X, but the court then rejected linguistic interpretation and decided for X, we annotated only the historical interpretation.
3.3 Holistically labelling judgments as formalistic
Although arguments are crucial to determine prevailing reasoning style, formalism is not just about quantity of arguments. In our pilot studies, simply counting argument frequencies proved important, but insufficient. Some decisions dominated by formalistic arguments turned out non‑formalistic once assessed holistically—against the core tenets, the whole decision, its context, facts, and procedural history. Although quantification of arguments seems to be crucial prerequisite to study formalism and would likely suffice in most cases as our quantitative results show, it can also be misleading in some cases. There are three main reasons:
First, arguments differ in weight and form. A judgment may invoke a fundamental principle once or twice in a decisive paragraph, yet be packed with dozens of formalistic arguments in peripheral sections on procedure. Such a decision likely would be, in substance, non‑formalistic, but simple counting would suggest the opposite. Argument clarity also varies: some arguments are explicit; others are barely there and must be inferred from context. The problem deepens when courts cite statutory provisions that themselves embed non‑formalistic considerations (e.g., principles), blurring whether the reasoning is formalistic or not. And some “principles” are formalistic by design—think of the ”concentration principle,” which strictly bars new evidence in later stages of the proceedings. While arguments significantly matter, their quantification alone might therefore mislead.
Second, context matters. A Supreme Court decision critising lower courts for excessive formalism may still be filled with citations to statutory text and case law (formalist arguments) while ultimately pushing for a purposive, non-formalistic decision-making.
Third, the last two tenets, i.e., dismissal on formal grounds and insufficient reasoning, are not capturable by simply counting arguments.
To address this limitation, we annotated each decision with a binary label ‘formalistic/non-formalistic’. The expert decision is based on an holistic assessment of the entire case. The evaluations is complex and considers multiple contextual factors beyond argument counting: adherence to the five core tenets of formalism, the overall approach to legal interpretation, the weight and centrality of different arguments, whether the court critiques previous courts for formalistic reasoning, and the broader procedural and substantive context of the case.
3.4 Annotation process
The systematic annotation of Czech Supreme Courts’ decisions was conducted between July and September 2024, building on a pilot study from early 2024 that analyzed 160 decisions and helped refine the methodology. We employed four law students from the Faculty of Law, Charles University, as research assistants, that were lead by a PhD student in legal argumentation. All of them had a background in legal theory and argumentation.
The process had four phases: Introduction (Week 1), Training (Weeks 2–6, 60+ decisions), Coding (Weeks 7–12, 272 decisions with regular reviews), and Finalization (Week 13). The team met on weekly bases to discuss the annotations in the Introduction and Training phase. We used the INCEpTION [tubiblio106270] interface for accessible annotation. We followed methodological guidance for the legal domain and drawed mainly from content analysis literature [hall2008, Ovádek_Schroeder_Zglinski_2024] and recommended practices by braun_i_2023-1 for annotating legal NLP datasets. This approach involved several hundred hours of training and guideline development, with at least two independent annotators coding each decision in the MADON dataset. We measured inter-annotator agreement and report on the annotation process. In total, the annotation took 1,000+ hours, covered more than 450 decisions including pilot studies (272 after pilot), and produced the MADON dataset with 272 decisions consisting of 9,183 annotated paragraphs, including 1,913 legal arguments in total and 1,237 paragraphs with at least one argument.
3.4.1 Inter-annotator agreement and dataset curation
We measured inter-annotator agreement to asses reliability of our annotations. We calculated two different inter-annotator agreement metrics corresponding to our two types of labels: Cohen’s kappa for the holistic label and Krippendorff’s unitized Alpha for argument types.
Our holistic assessment achieved a Cohen’s kappa coefficient of 0.65 on the overall category formalistic/non-formalistic, indicating “substantial agreement” among annotators despite the complexity of the assessment [Landis_Koch_1977]. This success yields two significant findings for future empirical research. First, research assistants can be effectively trained to achieve sufficient inter-coder agreement on complex holistic assessments of formalistic/non-formalistic reasoning. Second, we found a strong correlation between decisions with higher frequencies of non-formalistic arguments and holistic nonformalistic assessments, suggesting that while the dual approach provides valuable validation, the relationship between argument types and overall formalism remains strong.
Table 1: Inter-annotator agreement in Krippendorff’s $\alpha_{u}$ for each argument type
Argument Part I Part II
Linguistic Interpretation (LIN) 0.54 0.28
Systemic Interpretation (SI) 0.35 0.42
Case Law (CL) 0.95 0.94
Doctrine (D) 0.94 0.90
Historical Interpretation (HI) 0.68 0.80
Principles of Law and Values (PL) 0.76 0.65
Teleological Interpretation (TI) 0.63 0.65
Practical Consequences (PC) 0.20 0.21
For argument type annotation, we calculated Krippendorff’s unitized Alpha across two parts of the annotation process (in total 272 decisions), see Table 1. According to Krippendorff, sufficient agreement requires $\alpha\geq 0.800$, while data with agreement $0.800>\alpha\geq 0.67$ can be used to draw tentative conclusions [Krippendorff_2019]. Half of our categories meet the 0.67 threshold, while three categories (Linguistic Interpretation, Systematic Interpretation, and Practical Consequences) demonstrate significantly lower reliability. This pattern reflects the inherent complexity of legal document annotation. Annotation of legal documents is generally considered challenging, and inter-annotator agreement is often lower [braun_i_2023-1, vanDijck2024].
Disagreements were resolved through “arbiter review” as suggested by [braun_i_2023-1] or a combination of arbiter review and “forced agreement” for complex cases – mainly the holistic label. The combination of arbiter review and forced agreement meant that annotators were required to discuss their disagreements and propose potential solutions, accompanied by justifications for their choices. The arbiter then reviewed both the suggested solutions and their justifications; when annotators agreed on how to solve initial disagreement but the arbiter disagreed with their consensus, the arbiter assigned the final label (i.e., the initial label of one of the annotators).
4 MADON dataset and first empirical findings
The MADON [^2] MADON stands for Mining Arguments to Debunk Old Narratives dataset comprises 272 legal decisions from Supreme Court and Supreme Administrative Court, the only two Supreme Courts. [^3] Like Germany, Austria, Poland and many other countries in Europe, Czechia has a separate Constitutional Court “above” the Supreme Courts. It hears mainly constitutional complaints about alleged violations of fundamental rights. The Constitutional Courts is significantly distinct from the Supreme Court and the Supreme Administrative Court and is not included in our dataset.
4.1 Stratified sampling of documents for expert annotation
The MADON dataset spans the period from 1997 to 2024, with a focus on decisions made between 2003 and 2023. To ensure a representative sample, stratified sampling was employed, balancing cases across time periods, court agendas (civil, criminal, and administrative), and types of cases (procedural and on the merits). This method ensured proportional representation across the Supreme Court’s civil branch (122 cases), criminal branch (58 cases), and the Supreme Administrative Court (90 cases). Within these subpopulations, we drew random samples.
4.2 MADON statistics
(a) Document length distribution
(b) Number of arguments distribution
Figure 1: Documents in MADON are (a) long on average but rarely exceeding 6,000 tokens, and contain 7 arguments on average (b) with some outliers with almost 80 arguments.
The MADON dataset contains 9,183 paragraphs with 1,913 annotated arguments. 765 paragraphs contain one argument, 472 paragraphs contain 2 or more arguments. 7,946 paragraphs (87%) contain no arguments. Individual decisions are highly heterogeneous. The shortest decision runs to 275 tokens, the longest to 22,831, with a mean of 2,923 tokens. The average number of arguments is 7, with a maximum of 81 arguments. Notably, 43 documents contain no arguments at all; see Figure 1 for histogram plots. This heterogeneity is caused by the fact that we intentionally represented all types of decisions, including procedural dismissals that are often short.
Figure 2: Without distinguishing between the two courts, the most prevalent argument type for the entire period 2003–2023 is Case Law (CL).
Argument types
The distribution of arguments reveals interesting patterns (see Figure 2). Linguistic interpretation, despite its theoretical importance for formalism, appears only 108 times in the whole MADON dataset. Another surprising aspect is the extremely high frequency of Case law (707 arguments) which we discuss below.
Table 2: Distribution of formalistic arguments through paragraphs and documents
Arguments
Linguistic
Interpretation
Systemic
Interpretation
Case Law
Doctrine
Paragraphs (frequency) 108 208 707 68
Documents (existence) 63 78 181 39
Table 3: Distribution of non-formalistic arguments through paragraphs and documents
Arguments
Historical
Interpretation
Principles of Law
and Values
Teleological
Interpretation
Practical
Consequences
Paragraphs (frequency) 58 311 315 138
Documents (existence) 22 117 124 68
Holistic label
The distribution of holistic labels differs across courts. The SC had 116 formalistic (64%) and 66 non-formalistic (36%) decisions for a total of 182 cases. The SAC had 45 formalistic (50%) and 45 non-formalistic (50%) decisions for a total of 90 cases. Combined, there were 161 formalistic (59%) and 111 non-formalistic (41%) decisions out of 272 total cases.
4.3 Key legal empirical findings from the MADON dataset
The MADON dataset allows us to get the first systematic picture of both Czech Supreme Courts’ reasoning practices. Contrary to the Tale of Two Courts discussed in Section 2.2, i.e., a Tale of formalistic SC and non-formalistic SAC, we found that during the SAC’s initial decade (2003-2013), both courts exhibited remarkably similar levels of formalism on most indicators. For instance, both courts had approximately two-thirds of their decisions classified as formalistic, with the SAC showing even slightly more formalism on some indicators.
However, the temporal trends in Figure 3 show a notable divergence which started around 2011 and continued throughout the second decade (2014-2023), where the SAC shifted towards non-formalistic reasoning. During this second period, the proportion of non-formalistic SAC decisions increased by 70 percent (Figure 3c), and the average number of non-formalistic arguments per decision rose by more than 100 percent (Figure 3b), while the SC’s practices remained relatively stable. This divergence is also evident in the ratio of non-formalistic to formalistic arguments (Figure 3a) and the decreasing share of SAC decisions with zero non-formalistic arguments (Figure 3d). This indicates that the Tale of Two Courts—the narrative about formalistic SC and non-formalistic SAC—became close to reality only in the later period, if ever.
(a) Non-formalistic/formalistic argument ratio
(b) Average non-formalistic arguments per decision
(c) Share of non-formalistic decisions
(d) Share of decisions with zero non-formalistic arguments
Figure 3: Temporal trends in judicial reasoning patterns: Supreme Court (SC) vs. Supreme Administrative Court (SAC), 2003-2023
Furthermore, the findings challenge stereotypes about CEE legal reasoning. As can be seen from Figure 2, both the SC and SAC use text-based arguments scarcely: linguistic interpretation (LIN) accounts for 5.7% of all arguments across both courts in 2003–2023. By contrast, non-formalistic arguments from legal principles and values (PL) and teleological interpretation (TI) are much more common, together making up 32.3% of all arguments. Case law (CL) is the most common type at 37.4%. Notably, case law was by far most dominant argument type already in 2006 at 29%; this indicates that while its use has increased over time (both in percentage share and absolute numbers), it was already the most frequent category of legal arguments at that earlier point in the study period.
These results demonstrate that Czech judicial reasoning, particularly its heavy reliance on case law and non-formalistic argumentation, diverges from traditional definitions of formalism prevalent in both CEE and U.S. contexts, where formalism is connected to textualism. We thus empirically refute the notion that Czech courts are text-bound and had been precedent-averse [manko_weeds_2013, matczak_eu_2015, Spamann_2024].
4.4 MADON dataset for NLP experiments
For the experiments described in the upcoming section, we split the MADON dataset into train (70%), validation (20%), and test (10%) sets. The split was stratified to proportionally represent holistic labels and issuing court throughout the subsets, as summarized in Table 4.
Table 4: MADON dataset splits with label and court distributions
Train Validation Test Total
Holistic Labels
   Formalistic 112 (59.3%) 32 (59.3%) 17 (58.6%) 161 (59.2%)
   Non-Formalistic 77 (40.7%) 22 (40.7%) 12 (41.4%) 111 (40.8%)
Court Distribution
   Supreme Court 127 (67.2%) 36 (66.7%) 19 (65.5%) 182 (66.9%)
   Supreme Admin. Court 62 (32.8%) 18 (33.3%) 10 (34.5%) 90 (33.1%)
Total 189 54 29 272
5 Experiments
5.1 Experiment setup and tasks
We design three different experiments to investigate the connection between court document texts, arguments, and formalism in the MADON dataset. In general, mining arguments is often split into several subtasks [eger-etal-2017-neural] and we follow this pattern by defining our subtasks into the following: As an initial probe whether models are capable of detecting arguments at all, we train it to predict the presence and absence of arguments in a paragraph. We do not distinguish between a single argument and multiple arguments and focus only on detecting the presence. This experiment design also serves as a sanity check to see if our models are able to distinguish argumentative and non-argumentative paragraphs. The second and most difficult task requires the extraction of the exact argument types from a given paragraph. Since there can be more than one argument in a paragraph, we formulate the problem as a multilabel classification problem with a binary representation for the presence of each argument type separately. If there is no argument present, the output should be ‘False’ for each argument type. Finally, we tackle the question of formalism for the entire document. This third task should also provide us with insights whether models are capable of understanding the document holistically without having additional labels. Here we experiment with different approaches and pipelines; the simplest approach passes the entire text of the decision to the model and performs a binary classification (formalistic/non-formalistic). In the following, we refer to our experiments as:
Task 1: Argument presence detection
We classify paragraphs into argumentative and non-argumentative based on the presence or absence of any argument in it.
Task 2: Argument type classification
Entire paragraphs are classified in a multilabel setting into the eight types of traditional legal argument types described in Section 3.2.
Task 3: Holistic formalism classification
We classify entire judicial decisions as formalistic or non-formalistic.
5.2 Models
For all three tasks, we use large language models (LLMs) that can handle large documents efficiently. However we also keep computational efficiency at inference time in mind to enable empirical analysis of the entire judiciary of 230,000 Supreme Courts decision in future work. We therefore experiment with two complementary state-of-the-art architectures: the decoder-only Llama 3.1 8B [ai2024llama3] and the encoder-only ModernBERT [modernbert]. Llama 3.1 is trained on multilingual data, making it suitable for our Czech data, while ModernBERT provides a strong encoder baseline with long-context support. Moreover, for the holistic formalism classification task, we also experiment with feature-based multi-layer perceptron (MLP).
5.2.1 Llama
We use two variants of Llama 3.1 8B: the base model without instruction fine-tuning, which is suited for multilabel task finetuning, and a base model version which received continued pretraining (CPT) on Czech legal texts (Section 5.2.3 below). We also test the performance of instruction tuning (IT) on Llama 3.1 70B with different variants of prompting using annotation guidelines in zero-shot and few-shot settings. To optimize computational costs, we also test the performance of parameter-efficient fine-tuning (PEFT, via LoRA adapters [hu2021lora]). However, we find that instruction tuning performed significantly worse than fully fine-tuned models.
5.2.2 ModernBERT
ModernBERT 395M [modernbert] is a recent encoder-only transformer model supporting up to 8,192 tokens. Compared to earlier models of the BERT family, it integrates modifications such as Rotary Positional Embeddings (RoPE) [su-etal-2023-roformer], GeGLU activations, alternating local–global attention, and Flash Attention [dao-etal-2022-flashattention]. These features make it a suitable encoder baseline for our classification tasks, overcoming the token limitation and English bias of Legal-BERT [chalkidis2020legalbert].
5.2.3 Continued pretraining of transformers
Neither Llama 3.1 8B nor ModernBERT was originally trained for the legal domain or the Czech language specifically. To address this limitation, we perform continued pretraining of both models using Czech legal data. This approach allows us to investigate whether domain- and language-specific adaptation improves performance, building on previous findings in legal NLP [zheng2021whendoes, Habernal.et.al.2023.AILaw]. For pretraining, we curate a corpus of approximately 300,000 legal decisions from the Supreme Court (SC), Supreme Administrative Court (SAC), and regional administrative courts, excluding all decisions already included in the MADON dataset to prevent data contamination. [^4] Data available at https://github.com/trusthlt/madon/.
Due to computational constraints, we continue pretraining the Llama 3.1 8B base model for one epoch, using an effective batch size of 16 and a maximum input length of 32,000 tokens, which is large enough to fully cover most of the documents. We use unsloth [^5] https://unsloth.ai to optimize the training process.
We follow a two-phase pretraining process for ModernBERT. First, we train a custom tokenizer using byte-pair encoding on our pretraining dataset. We then continue pretraining from the last checkpoint of ModernBERT-large using a masked language modeling objective.
In phase 1, we include documents with fewer than 8,192 tokens (the model’s input limit), train with a 30% masking probability and an effective batch size of 32. In phase 2, we reduce the effective batch size to 8 and the masking probability to 15% and train using the remaining documents.
5.2.4 Multi-layer perceptron
For holistic formalism classification, we also experiment with an alternative input representation. We simply use the annotated argument types and extract various statistics into a document-level feature vector which we feed into a multi-layer perceptron (MLP; 2 hidden layers with 20 and 50 neurons, respectively). [^6] Simpler linear models such as logistic regression did not produce any usable results. The feature set extracted from a document includes the document length in tokens, number of annotated arguments, average argument length, and for each argument type from our inventory its relative frequency in percent. Note that this feature-based approach is only applicable to documents annotated with arguments (which is the case for the MADON dataset) but cannot be used for inference on purely textual data, unlike the transformer models. We will touch on mitigating this limitations later in Section 5.3.3. For further training details see Appendix B.
5.3 Results
For all three tasks, we rely on the macro-averaged F1 score. To provide a balanced evaluation, we distinguish between two complementary perspectives: the positive F1 score ($\text{F1}^{+}$), which measures how well the model identifies labels that are actually present, and the negative F1 score ($\text{F1}^{-}$), which measures how well the model avoids predicting labels that are absent.
For Task 1 (argument presence detection) and Task 3 (holistic formalism classification), the evaluation is binary: either an argument (resp. formalism) is present or not. We compute the macro-averaged F1 score over the two classes (present vs. absent), following the standard definition.
For Task 2 (argument type classification), the evaluation is multi-label with an inventory of eight argument types $L=\{\text{LIN},\text{SI},\text{CL},\text{D},\text{HI},\text{PL},\text{TI},\text{PC}\}$. Initial experiments with standard metrics for multi-label evaluation, such as Hamming distance, did not capture the nuances in misclassification and did not help us interpret the results across various argument types, which is essential for empirical legal analysis. We therefore propose an evaluation metrics tailored to our task.
Each label $\ell\in L$ is treated as an independent binary classification problem. Concretely, for a given label $\ell$:
  • if $\ell$ is annotated in the gold data and predicted, this counts as a true positive (tp), contributing to $\text{F1}^{+}$,
  • if $\ell$ is not annotated but predicted, this is a false positive (fp), which lowers both $\text{F1}^{+}$ and $\text{F1}^{-}$,
  • if $\ell$ is annotated but not predicted, this is a false negative (fn), which also lowers both $\text{F1}^{+}$ and $\text{F1}^{-}$,
  • if $\ell$ is neither annotated nor predicted, this is a true negative (tn), contributing to $\text{F1}^{-}$.
For each label $\ell$, we compute both the positive and negative F1 scores, namely
$\text{F1}^{+}_{\ell}=\frac{2\cdot\mathrm{tp}_{\ell}}{2\cdot\mathrm{tp}_{\ell}+\mathrm{fp}_{\ell}+\mathrm{fn}_{\ell}},\qquad\text{F1}^{-}_{\ell}=\frac{2\cdot\mathrm{tn}_{\ell}}{2\cdot\mathrm{tn}_{\ell}+\mathrm{fp}_{\ell}+\mathrm{fn}_{\ell}}.$ (1)
We then define the macro-F1 for a given label $\ell$ as
$\text{Macro-F1}_{\ell}=\frac{\text{F1}^{+}_{\ell}+\text{F1}^{-}_{\ell}}{2}.$ (2)
Finally, the overall macro-F1 is obtained by averaging across all labels:
$\text{Macro-F1}_{\text{all}}=\frac{1}{|L|}\sum_{\ell\in L}\text{Macro-F1}_{\ell}.$ (3)
5.3.1 Results for Task 1: Argument presence detection
The top-performing models reach macro F1 scores above 82%, with macro recall exceeding 79% for the best ModernBERT configurations. For this binary classification task, ModernBERT and Llama achieve very similar results at their best configurations. The highest macro F1 scores are 79.5% for Llama 3.1 8B Baseand 82.5% for ModernBERT Large, with the CPT variants performing within a small margin of these values, as shown in Table 5.
Table 5: Result of argument presence detection (Task 1). Best macro F1 in bold.
Model Macro Precision Macro Recall Macro F1
Majority baseline 43.6 50.0 46.6
Random baseline 50.3 50.7 42.0
Llama 3.1 8B Base 87.1 75.2 79.5
Llama 3.1 8B Base + PEFT 53.0 55.1 49.7
Llama-3.1-8B-Czech-Legal 85.8 65.9 70.7
Llama-3.1-8B-Czech-Legal + PEFT 53.7 55.7 51.3
ModernBERT Large 86.2 79.8 82.5
ModernBERT-large-Czech-Legal 86.0 80.1 82.6
Given that ModernBERT is a much smaller model compared to Llama 3.1 8B, its ability to slightly outperform the larger Llama model suggests that this task does not require the full capacity or extended context window of a large-scale LLM. This opens the possibility of deploying smaller, more efficient models in practice without sacrificing accuracy. Continued pre-training on Czech legal texts did not enhance performance for either architecture. In fact, for Llama, the CPT variant underperforms the base model (70.7% vs. 79.5% macro F1), and for ModernBERT the difference is negligible (82.6% vs. 82.5% macro F1). This suggests that the task can be solved effectively using base models.
Surprisingly, parameter-efficient fine-tuning via LoRA led to a severe drop in performance. Llama 3.1 8B Base + PEFT configuration fell from 79.5% to 49.7%, while the Llama-3.1-8B-Czech-Legal + PEFT variant performed similarly poorly at 51.3%. This indicates that for argument presence detection, full fine-tuning is essential to achieve high performance.
5.3.2 Results for Task 2: Argument type classification
Table 6 presents per-label metrics and macro-averaged results for Task 2. By macro-average F1, the strongest performance is achieved by the fully fine-tuned Llama 3.1 8B Basewith Asymmetric Loss (Base + Asy), reaching a macro F1 of 77.5, an average across all 8 labels.
This configuration consistently outperforms all baselines and other model setups, indicating that full fine-tuning allows the model to better capture the distribution and complexity of argument types in the MADON dataset. ModernBERT achieves competitive results, with its best performance coming from the CPT + Asy variant (macro F1 of 71.6%), though still notably behind the fully fine-tuned Llama. In contrast, PEFT-based Llama 3.1 8B configurations yield much lower macro F1 scores (49.6% at best), performing only slightly above the majority baseline (49.4%) and suggesting that the parameter-efficient setup, as implemented here, is insufficient for this task.
Across all architectures, replacing Binary Cross-Entropy with Asymmetric Loss consistently boosts macro-average performance. The gain is most evident in ModernBERT (Base: 62.2% $\rightarrow$ 67.3%; CPT: 62.9% $\rightarrow$ 71.6%) and in the fully fine-tuned Llama 3.1 8B (74.6% $\rightarrow$ 77.5%). This pattern aligns with the class imbalance observed in the MADON dataset, where the negative class dominates across all labels, and Asymmetric Loss better emphasizes minority positive instances.
The effect of continued pre-training (CPT) varies by architecture: for ModernBERT, ModernBERT-large-Czech-Legal leads to substantial gains, while for Llama 3.1 8B full fine-tuning, Llama-3.1-8B-Czech-Legal slightly underperforms the Base variant.
Table 6: Task 2: Per-label metrics and macro average. BCE = Binary Cross-Entropy, Asy = Asymmetric Loss, PEFT = Parameter-Efficient Fine-Tuning, CPT = Continued Pre-Trained Model. Bold means best macro-averaged F1.
Model Class LIN SI CL D HI PL TI PC Mac-Avg
Baselines Majority Pos 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Neg 99.4 98.8 96.6 99.9 99.6 97.7 98.3 99.2 98.7
Avg 49.7 49.4 48.3 49.9 49.8 48.9 49.2 49.6 49.4
Random Pos 2.2 4.4 12.7 0.8 1.5 7.4 6.1 1.8 4.6
Neg 66.5 64.2 66.9 68.0 67.2 67.1 65.2 65.8 66.3
Avg 34.3 34.3 39.8 34.4 34.3 37.2 35.7 33.8 35.5
Llama 3.1 8B (PEFT) Base + BCE Pos 0.8 4.3 17.6 0.0 3.4 6.1 4.4 2.7 4.9
Neg 94.2 92.1 80.6 92.3 93.3 79.2 84.7 88.1 88.1
Avg 47.5 48.2 49.1 46.1 48.4 42.7 44.5 45.4 46.5
Base + Asy Pos 0.0 1.6 18.9 0.0 0.9 3.3 1.5 3.1 3.7
Neg 97.5 97.4 90.8 96.4 97.3 94.0 94.6 96.4 95.6
Avg 48.8 49.5 54.9 48.2 49.1 48.6 48.1 49.8 49.6
CPT + BCE Pos 0.0 6.2 17.4 0.0 2.3 7.2 2.1 0.7 4.5
Neg 97.1 93.6 82.2 91.1 92.2 83.5 90.0 95.7 90.7
Avg 48.5 49.9 49.8 45.6 47.2 45.4 46.0 48.2 47.6
CPT + Asy Pos 0.0 1.1 13.7 0.0 2.4 2.3 1.6 1.7 2.9
Neg 98.7 97.5 90.9 97.8 98.2 92.9 96.2 98.4 96.3
Avg 49.3 49.3 52.3 48.9 50.3 47.6 48.9 50.0 49.6
Llama 3.1 8B (Full) Base + BCE Pos 41.4 30.3 81.7 75.0 14.1 75.4 48.2 34.7 50.1
Neg 99.5 98.9 98.8 99.9 99.6 98.9 98.6 99.4 99.2
Avg 70.4 64.6 90.2 87.5 56.9 87.1 73.4 67.0 74.6
Base + Asy Pos 57.6 34.8 79.5 63.9 36.2 74.8 55.1 46.9 56.1
Neg 99.4 98.3 98.5 99.8 99.6 98.7 98.2 99.1 99.0
Avg 78.5 66.5 89.0 81.9 67.9 86.8 76.7 73.0 77.5
CPT + BCE Pos 36.8 16.6 81.0 75.0 0.0 75.8 51.4 26.2 45.3
Neg 99.5 98.8 98.7 99.9 99.6 98.9 98.7 99.3 99.2
Avg 68.1 57.7 89.9 87.5 49.8 87.4 75.0 62.8 72.3
CPT + Asy Pos 56.3 37.0 78.2 67.2 19.4 70.7 57.6 43.0 53.7
Neg 99.4 98.4 98.3 99.9 99.5 98.5 98.3 99.0 98.9
Avg 77.8 67.7 88.2 83.5 59.4 84.6 77.9 71.0 76.3
ModernBERT Base + BCE Pos 9.6 5.9 69.7 24.4 9.3 49.1 27.7 10.0 25.7
Neg 99.2 98.5 98.0 99.8 99.4 98.1 98.0 98.9 98.8
Avg 54.4 52.2 83.9 62.1 54.3 73.6 62.9 54.4 62.2
Base + Asy Pos 19.4 17.8 67.5 67.2 23.3 48.5 32.2 11.4 35.9
Neg 99.3 98.3 97.5 99.9 99.6 97.8 97.7 99.0 98.6
Avg 59.3 58.0 82.5 83.5 61.5 73.1 65.0 55.2 67.3
CPT + BCE Pos 9.3 13.3 67.4 28.6 3.4 49.2 37.4 12.7 27.7
Neg 97.0 97.8 97.8 98.8 98.7 98.2 97.7 98.3 98.0
Avg 53.2 55.5 82.6 63.7 51.1 73.7 67.5 55.5 62.9
CPT + Asy Pos 29.9 32.7 75.1 59.3 21.9 63.2 39.5 33.6 44.4
Neg 99.1 98.4 98.2 99.8 99.3 98.1 98.0 99.2 98.8
Avg 64.5 65.6 86.7 79.5 60.6 80.6 68.7 66.4 71.6
5.3.3 Results for Task 3: Holistic formalism classification
Table 7: Results of holistic formalism classification (Task 3). Bold = highest macro F1 (per block).
Model Macro Precision Macro Recall Macro F1
Baselines
Majority baseline 29.3 50.0 36.9
Random baseline 42.7 42.7 41.4
Using gold-annotated arguments as features
MLP 91.9 91.5 91.7
End-to-end models
Llama 3.1 8B Base 77.1 75.5 75.9
Llama 3.1 8B Base + PEFT 72.5 60.5 57.6
Llama-3.1-8B-Czech-Legal 79.9 77.5 78.0
Llama-3.1-8B-Czech-Legal + PEFT 79.6 62.9 60.3
ModernBERT Large 74.1 73.9 73.8
ModernBERT-large-Czech-Legal 82.4 82.0 82.2
Multi-step pipelines
Basic 84.0 86.3 83.2
With Filtering 83.8 87.2 82.8
Gold labeled arguments as features
Arguments are crucial for holistic evaluation of formalism. Using human expert annotations of arguments from MADON dataset as features, a MLP reaches macro-F1 of 91.7% in classifying decisions, substantially higher than any contemporary end-to-end LLMs (see Table 7). The comparison is obviously not fair, as the input of LLMs is the entire decision plain text while the MLP knows the true arguments, but it shows that the presence, absence, and distribution of specific argument types do capture the essence of judicial formalism, even though the legal experts considered other aspects (e.g., context of the case) when assigning the holistic label for the MADON dataset.
(a) Non-formalistic documents
(b) Formalistic documents
Figure 4: Summary plot for SHAP values of all test examples
Not all arguments contribute equally. Figure 4 summarizes an explainability method using SHAP analysis [lundberg2017unified] and shows that Principles of Law and Values (PL) arguments are the strongest predictor of non-formalism: higher frequencies of PL are closely linked to non-formalistic decisions. Teleological Interpretation (TI) and the total number of arguments also correlate positively with non-formalism, which aligns with our theoretical definition that describes formalism as the avoidance of purposive reasoning and limited number of arguments. By contrast, Systematic Interpretation (SI), though classified as formalistic in theory, has little predictive value.
Contemporary end-to-end LLMs for holistic classification
Direct classification of decisions as non/formalistic—without argument features—gives very promising results. Continued pretraining improves both model families, and ModernBERT-large-Czech-Legal achieves 82.2% macro-F1, the best among end-to-end models. ModernBERT outperforms Llama even though Llama can process the entire document context without a cut-off due to a token limit. This indicates that lexical and semantic cues, which also play an important role in expert annotation, may be more valuable than full-context processing. ModernBERT outperfroms the baselines by 40+ points, yet still falls behind the argument-based approach by about ten points.
Multi-step pipeline for holistic classification
We experimented with different approach to holistic classification. Since even the best LLMs fall behind the feature-based performance, we designed a multi-step pipeline that uses features and mirrors how legal scholars would evaluate decision and assess formalism: first filtering irrelevant passages, then identifying argument types, and finally assigning a holistic label formalistic/non-formalistic. This architecture provides better explainability and even slightly improves performance. The pipeline consists of three stages:
  1. 1.
    In the first stage, ModernBERT Large identifies which paragraphs contain legal arguments, filtering out ca 85% of the decisions’ text (i.e. non-argumentative parts of the decisions; these are often repetitive or procedural texts usually not crucial for assessing formalism).
  2. 2.
    In the second stage, Llama 3.1 8B Base (+ Asy), classifies the filtered text into argument types such as Teleological Interpretation, Principles of Law, and Systematic Interpretation.
  3. 3.
    Finally, MLP extracts features from the identified arguments to make the final formalistic/non-formalistic classification, using our previous finding that argument frequency strongly predicts the overall non/formalistic nature of the decision. See Figure 5 for a visual depiction.
Figure 5: A multi-step pipeline combining several distinct customized LLMs and MLP delivers strong and explainable performance for holistic formalism classification.
The three stage pipeline yields 82.8% macro F1 and slightly outperforms the best end-to-end model. Interestingly, if we omit the filtering stage and apply Llama to entire decisions, the score rises marginally to 83.2%. These results provide two key findings. First, decomposing NLP classification task according to legal theory does not sacrifice performance; on the contrary, the pipeline scores slightly better than end-to-end models and provides greater explainability (due to the reliance on features in MLP that can be explained by SHAP methods). Second, filtering as much as 85% of the text of the classified documents has negligible impact on performance for this task but significantly safes compute costs. Overall, it proved beneficial to let legal theory to guide the NLP architecture for this classification task and opt for feature based approach.
To sum up, our results show that computational methods can classify complex judicial philosophies even in non-English judicial decisions.
6 Discussion
Our study provides three main findings. First, we developed state-of-the-art NLP methods that are able to detect argumentative paragraphs, can classify traditional types of legal arguments but struggle with some of them, and can identify formalistic versus non-formalistic decisions in Czech Supreme Courts judgments with high accuracy. Second, we show that decomposing legal classification tasks and using more explainable, segmented pipelines across model families can improve performance and explainability while reducing compute costs. Third, the results refute dominant narratives about judicial formalism in Central and Eastern Europe, particularly the Tale of Two Courts in Czechia. We believe that achieving very promising judicial philosophy classification in a non-English setting represents a significant advance for legal NLP and the rising field of computational legal studies.
6.1 Contributions to legal NLP
This is the first LLM-based argument mining study on CEE courts. Our experiments and lessons learned on non-English court judgments could benefit the legal NLP community. Parameter-efficient fine-tuning was inadequate, whereas continued pretraining proved important for ModernBERT (namely for argument-type classification and holistic formalism classification). By contrast, Llama benefited primarily from full fine-tuning rather than continued pretraining. Class imbalance—prevalent in legal datasets—needed explicit treatment: for multi-label argument-type classification, Asymmetric Loss consistently outperformed Binary Cross-Entropy across model architectures, improving macro-F1 by up to nine points. Finally, macro-averaged metrics were essential for multi-label evaluation; the scores would be otherwise dominated by the majority negative class for each argument type.
We also show that interdisciplinary combination of legal theory and NLP can be fruitful. Rather than classifying reasoning style at the decision level directly, this NLP task benefits from decomposing the pipeline according to legal theory: first, filter legally relevant parts of the text, then classify traditional legal arguments, and only then classify legal formalism. We expect that decomposing classification tasks in accordance with how lawyers or legal theorists approach problems could boost explainability and performance in other legal NLP tasks.
Crucially, our pipeline to investigate reasoning patterns in court decisions is replicable in other jurisdictions (e.g., Poland, Germany, US) and paves the way for comparative legal NLP studies.
6.2 Implications for computational legal studies
A key strand of computational legal studies is “law as data”—an empirical approach to legal research that uses computational tools to study courts and judicial culture [Chau_Livermore_2024]. We show how LAM advances this agenda: imperfect argument mining enables very promising classification of judicial philosophies. Even when argument detection is noisy for some categories, the resulting distributions can capture the key aspects of judicial reasoning we were interested in.
More broadly, our work illustrates how computational legal studies using LAM can move legal research beyond anecdotes. Empirical analysis at scale can challenge entrenched narratives, provide evidence on judicial practice, and support the legitimacy of courts by addressing criticisms with systematic data. The release of our MADON dataset (with an accompanying raw corpus of 300k decisions) and all models provides the necessary resources for this research agenda.
6.3 Implications for legal scholarship
Our findings challenge decades of assumptions about CEE courts. The divergence between the Supreme Court and Supreme Administrative Court appears only after 2011, and seem not be a direct consequence of communist-era legacies as commonly claimed [kuhn_judiciary_2011-2, matczak_constitutions_2010-2]. Both courts use textual arguments scarcely (5.7% of all arguments) while case law dominates (37.4%). This contradicts the claims that CEE courts are text-bound and had been precedent-averse [manko_weeds_2013, bystranowski_formalist_2022-1, Spamann_2024]. This preliminary empirical findings await further verification with a large scale analysis.
6.4 Implications for legal practice
For legal practice, LAM enables efficient document analysis. Paragraph-level filtering achieves 82.6% macro-F1 in identifying argumentative content, letting lawyers focus on argumentative passages which are often the crucial ones. Argument classification could help practitioners quickly find concrete argument types or specific reasoning patterns across single documents as well as large document collections (e.g., to assess what type of argumentation prevails in most filings and how they correlate with success rates at courts).
Beyond assessing argumentation patterns, LAM supports often very complicated evaluation of legal argumentation. Since each argument type has distinct evaluation criteria, automated detection simplifies the analysis. Practitioners can verify whether cited cases actually apply or whether teleological arguments connect to statutory purposes. Since evaluation is central to both practice and argument theory (lawyers want to know which arguments hold and which not), we believe future research could automate some parts of it. For instance, we could train models to check whether used teleological arguments indeed align with statutory purpose, a simple binary classification on top of detected arguments. This would create a two-stage method that would first identify arguments and then evaluate their strength.
Our approach also enables studying relationships between argumentation and case outcomes. Decisions with certain argument patterns may face higher reversal rates, which could inform both courts and participant (e.g. their litigation strategies). Courts and practitioners could also benefit from data about prevailing argument patterns of specific courts (and senates) for specific legal problems. For instance, if a lawyer knew she is preparing a filing for a court senate that predominantly relies on linguistic interpretation and rarely employs teleological arguments, she could strategically emphasize text based arguments to strengthen her position and improve her chances of success.
6.5 Limitations and future directions
Our research has several limitations.
First, empirical findings are restricted to 272 decisions, which, while representative through various sampling methods, limits generalizability. Nonetheless, this issue shall be solved by the large scale analysis with our models we plan to pursue as a next step.
Second, our formalism classification employs a CEE-specific definition that may be only partially applicable outside CEE. For instance, we classify Historical Interpretation (HI) as non-formalistic following CEE literature, whereas German or US scholarship might consider it formalistic. However, this limitation has minimal practical impact since HI arguments were extremely scarce in our dataset. Besides, our approach to formalism reflects our research objective: empirically testing existing claims about CEE formalism rather than proposing new theoretical framework. We conducted thorough literature review of prominent or recent CEE scholarship on the topic and extracted the core tenets of formalism to enable empirical verification of prevailing claims. This methodology aligns with our goal of replacing anecdotal claims with systematic evidence. And last but not least, the main core tenets seem to be shared outside CEE, especially the reliance on text-based arguments, focus on the most applicable rule while ignoring broader principles, and adherence to formal procedure to avoid substantive reasoning.
Third, the sparse occurrence of text-based arguments (5.7%) partially reflects our research design choices in our annotation guidelines. We required courts to engage at least partially with statutory language—analyzing wording, referencing dictionaries, or discussing grammar—rather than merely citing provisions. However, our threshold remained low: any reference to wording or explicit textual analysis qualified as linguistic interpretation. If we counted every single reference to a statute as linguistic interpretation, the frequency would be much higher; however, we believe our approach aligns better with legal argumentation theory, which treats analyzing wording, referencing dictionaries, and discussing grammar as the core aspects of linguistic interpretation.
Lastly, both empirical and NLP components focus exclusively on the two Czech apex courts. We cannot generalize empirical findings to lower courts or Constitutional Court nor simply use the models on judgments of these courts. Although applying the models especially to lower court decisions would be valuable to provide broader picture of the whole legal system, their applicability requires further examination and testing.
Several research directions follow. First, we plan to employ the trained models to the full corpus of 230,000 apex-court decisions to validate and refine our empirical results. Second, testing performance on lower-court decisions will show whether the approach needs adaptation across different courts in one jurisdiction. Third, extending our methodology and pipelines to other CEE jurisdictions would better test the existing claims about CEE region and would indicate whether the Czech patterns are country specific. Comparative LAM across CEE and non-CEE jurisdictions could directly test claims about CEE’s distinctiveness, or more broadly, verify existing claims about civil and common law reasoning that had been recently heavily critiqued [Spamann_2024]. Finally, we show that LAM could support related tasks in computational legal studies, including predicting outcomes from reasoning patterns. Whether it is indeed so, remains open.
7 Conclusion
We asked to which extent state-of-the-art NLP methods can detect and classify traditional types of legal arguments in non-english judicial decisions and whether they can distinguish between formalistic and non-formalistic reasoning styles at the decision level. After designing annotation guidelines for formalism and arguments and annotating 272 Czech Supreme Court decisions (9,183 paragraphs) resulting in the new MADON dataset, our extensive set of experiments shows that ModernBERT successfully distinguishes argumentative from non-argumentative paragraphs with a macro-F1 of 82.6%. Argument-type classification achieves macro-F1 of 77.5% using fully fine-tuned Llama 3.1 with asymmetric loss, though linguistic interpretation, systemic interpretation, and practical consequences remain hard to detect, unlike other argument types. Holistic formalism classification reaches impressive macro-F1 of 82.8% thanks to a novel three-stage pipeline, which at the same time also reduces computational costs and increases explainability. Empirically, both apex courts show similar formalism levels from 2003 to 2011, contrary to existing claims. After 2011, the SAC shifts toward non-formalistic reasoning while the SC remains stable. Text-based arguments are rare in both courts, while case law and purposive reasoning dominate.
Acknowledgements
This work has been supported by the German Research Foundation as part of the ECALP project (HA 8018/2-1), by the Research Center Trustworthy Data Science and Security (https://rc-trust.ai), one of the Research Alliance centers within the https://uaruhr.de, by the Charles University Grant Agency (project no. 185023), and by the Sylff Scholarship. We would like to thank Marlene Anzenberger, Michal Bobek, Zdenek Cervinek, Jonathan Choi, Corina Coupette, Dirk Hartung, Matthias Klatt, Pavel Ondrejek, Michael Preisig, Hubert Rottleuthner, Holger Spamann, Jed Stiglitz, and Barbara Zeller for their feedback on the early versions of this project, as well as all others who provided comments along the way. Special thanks to Václav Lipš, Vítek Eichler, Matyáš Barták, Marek Švajda for their tremendous research assistance.
Appendix A Arguments
For each argument type, our guidelines included a structured description that connects the theoretical concept to real‑world decisions. For most types, we (1) give a brief description, (2) list its subtypes, (3) note “trigger” phrases that usually signal its presence in a decision, (4) provide typical and borderline examples, (5) add comments and annotator tips and (6) include annotation chart.
All details are in the Guidelines. What follows is a brief description intended to illustrate how we understand the particular argument types:
Linguistic Interpretation (LIN)
This argument type focuses on the literal text of a legal provision. This includes analyzing the ordinary meaning of words, explicitly reflecting syntax or grammar rules, referencing explicit legal definitions within a statute, and employing reasoning a contrario. We considered trigger words like “unambiguously worded,” “diction,” or “formulation” as strong indicators.
Example:
“In addition to the above, the Supreme Administrative Court adds that the wording of Section 87e(1)(i)(1) of the Act on the Residence of Aliens is very unambiguous and leaves no room for a different interpretation.”
Systemic Interpretation (SI).
This argument involves interpreting a provision by placing it within the broader context of the legal system to ensure a ‘logical, consistent and non-contradictory interpretation’. This includes applying the collision rules (lex specialis, lex posterior, lex superior), narrowly construing exceptions, and explicitly performing conforming interpretation (either constitutional conforming or EU conforming interpretation) to align domestic law with higher-order legal norms.
Example:
“Since the provisions of Section 281 of the Criminal Procedure Code do not contain special provisions for decisions pursuant to Section 288(3) of the Criminal Procedure Code, the general provisions on the subject matter and local jurisdiction of the court (Sections 16 to 22 of the Criminal Procedure Code) apply.”
Case Law (CL).
This category captures express reliance on specific prior judicial decisions from domestic courts, the CJEU, or the ECHR. For an annotation to be made, the court must cite a particular case; a vague reference to “according to the case law” without a specific citation was not considered sufficient.
Example:
“In support of the conclusion, the Court also referres to the judgment of the Supreme Court of 11 May 2005, Case No. 30 Cdo 64/2004 (available at www.nsoud.cz).”
Doctrine (D).
This argument type refers to citations of academic or expert commentary, such as legal textbooks, scholarly articles, or commentaries. References to legal dictionaries are also marked as Doctrine, whereas non-legal dictionaries fall under Linguistic Interpretation.
Example:
“Current commentary literature on both the Code of Civil Procedure and the Civil Code derives from the nature of a power of attorney that the attorney‑in‑fact’s signature is not among its formal requirements (Smolík, P., § 28 [Granting a Power of Attorney and its Termination], in: Svoboda, K. et al., Občanský soudní řád, 3rd ed., Prague: C. H. Beck, 2021, pp. 121–122).”
Historical Interpretation (HI).
This argument type seeks to determine the lawmaker’s intent behind a provision. When court seeks to determine the intent, it typically refers to historical documents for legislative process, such as explanatory memoranda, stenographic records of parliamentary debates, or the circumstances of a law’s adoption. Due to its simmilarity, we also included arguments based on the assumption of a rational lawmaker.
Example:
““cf. the wording of (…) the special part of the explanatory report [“On points 4 and 5 (section 274)”] to the governmental bill No. 31/2019 Coll., which was debated by the Chamber of Deputies of the Parliament of the Czech Republic in its 8th term (2017-2021) as Document No. 71.”
Principles of Law and Values (PL).
This non-formalistic category includes appeals to higher-order norms that transcend the literal text of a single rule. It covers general legal principles (e.g., good faith, legitimate expectations, legal certainty) and explicit references to fundamental rights and values found in the Czech Constitution (incl. the European Convention on Human Rights) or EU Law. A principle does not need to be explicitly labeled as such by the court to be annotated.
Example:
“[…] which constitutes an interference with the right to a fair trial enshrined in Article 36(1) of the Charter”
Teleological Interpretation (TI).
This argument focuses on the purpose, goal, or “telos” of a legal provision, often signaled by terms like “purpose,” “aim,” “sense,” or “function”. This category also includes special forms of purpose-driven reasoning, such as arguments ad absurdum (to avoid a nonsensical result), a fortiori, and reasoning by analogy.
Example:
“An appeal is conceived as an extraordinary remedy and is therefore intended to remedy only serious legal defects in final decisions.”
Practical Consequences (PC).
This argument type involves the court considering the real-world effects and practical outcomes of its decision. Instead of focusing on the wording of legal norms, the court weighs what will happen if it chooses one interpretation over another, answering the question “What will it lead to?”. This can include impacts on the parties as well as the society.
Example:
““In the event of a contrary conclusion, it would be difficult to clearly identify and distinguish what personal manifestations of the taxi driver in the course of his interaction with the customer can still be characterized as his work activity and what already constitutes his private life.”
Appendix B Hyperparameters
MLP: We employ grid search hyperparameter tuning for our experiments. We use dropout for each hidden layers with the rates of 0.1. and 0.4 for the first and the second layer, respectively. We also apply early stopping with the patience of 3 epochs to prevent the overfitting. Finally, train the MLP model with the learning rate of 1e-3 and the batch size of 8. We also normalize the whole dataset based on the scaling factor of training dataset before the training, since the feature values are too sparse (see 4.2).
Appendix C Resources
Tables 8, 9 and 10 demonstrate the training and inference time for the best performing models for each task. Training means on the entire training set, inference means on the entire test set. Additionally, the utilized accelerators are shown in the last column.
Table 8: Runtime Task 1
Model Training Inference GPU
Llama (Base) 4h55m 43.0s A100(80GB)
ModernBert (CPT) 1h19m 3.8s RTX A6000 (48GB)
Table 9: Runtime Task 2
Model Training Inference GPU
Llama (Base + Asy) 6h36m 300.0s A100(80GB)
ModernBert (CPT + Asy) 2h12m 1.3s A100(80GB)
Table 10: Runtime Task 3
Model Training Inference GPU
Llama (CPT) 6h33m 13.0s A100(80GB)
ModernBert (CPT) 32m 1.3s RTX A6000 (48GB)